# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 8.3 Other
In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types. The First-Person Hand Action (FPHA) dataset (FPHA), published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants. It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction. Another notable example is the DP Dataset (DeePoint), released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition. Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances. It includes detailed annotations such as gesture timings and pointing directions, enabling both binary gesture classification (gesture vs. non-gesture) and directional estimation tasks. Finally, the SocialGesture (SocialGesture-CVPR2025-cao2025socialgesture) focuses on the recognition of natural gestures in multi-person environments. It comprises 9,889 in-the-wild videos containing 42,533 instances and provides fine-grained annotations including bounding boxes, relations (for example, a subject is pointing to a target) and Visual Question Answering (VQA) details.
@@ 8.4 Datasets Comparison
After examining the existing datasets for VHGR with significant impact on the current research, in this subsection we aim to determine pros and cons of the most important ones as well as the use cases that they fit. For gesture-based HCI, HaGRID (WACV_kapitanov2024hagrid) and its versions (HAGRIDv2_ARXIV) seem to be a reliable solution, containing an enormous number of samples corresponding to 19 distinct classes (that can be mapped to computer commands). Regarding that these gestures are also static, even simple architectures - like the ones discussed in sections 4 & 7 - can achieve sufficient performance. Furthermore, this dataset could also be used for pretraining, enhancing the backbone’s expression capability. However, its size implies that training takes a long time, which constitutes a deterrent in some cases, compared to other datasets for HCI like OUHANDS (ouhands). For gesture-based HRI, URGR (ULTRA-RANGE-UGV-Bamani) is the only one for UGV guidance and considers various recognition ranges, which may be crucial for real-world scenarios. Even though it contains a large number of samples, the small number of classes may limit the commands given by gestures and, therefore, the functionalities of the entire gesture-based interface. Regarding SLU datasets, for ASL recognition, OpenASL (OpenASL) is the largest one, including almost 100K RGB videos (288h in total) collecting from the web. MSASL (MS-ASL) comprises almost the same number of subjects, but it is much smaller. Furthermore, OpenASL was proposed for CSLR while for MSASL for IDSLR, thus it may be more suitable for real-world use cases, where the temporal boundaries are not known. BSL-1K (bsl-1K) is also a very large dataset, proposed for British IDSLR and comprising the largest number of samples (280K, 1060h in total). BOBSL (BOBSL-BBC) was proposed for British CSLR and comprises videos of 1467h in total making it the dataset with the longest duration. Phoenix (PHOENIX-2014) and Phoenix T (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) remain the most adequate dataset for German sign language, covering a large sign vocabulary. Nevertheless, the small number of samples and the low diversity - due to the limited number of signers and the constrained environment - make essential the collection of a new, larger and more diverse dataset for DGS. For Chinese SLR, NMFs-CSL (NMFs-CSL) is significantly smaller than CSLD (tim-slr_csld_dataseta), CSL (CSL500) and NCSL (2plus1-D-SLR_and_NCSL) but includes much more signs (1064 instead of less than 500). CSL-Daily (Back-Translation-CVPR) remains the only one for Chinese CSLR, covering a large sign vocabulary of 2000 signs in total.
@@ 9.2 Sequence-to-Sequence Metrics
For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used. Word Error Rate (WER) (VAC-CSLR; ADADOGLOU) is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
@@
Edit Distance Accuracy (FA-STGCN; CTCX_WANG2022123) is a complementary metric representing the ratio of correctly predicted elements in the sequence:
@@ 10.1 Gesture-Irrelevant Factors
The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios. Some of these factors include the following (EDenseNet; Mohammad_Mahmudul_static_hgr; 2DPSTPP; ULTRA-RANGE-UGV-Bamani; EnGesto; RAAR3DNet):
@@
Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference (FGDSNet; tuan_linh_pham_dang2023lightweight; DRCAM_TIM), object detectors for spatial normalization (ULTRA-RANGE-UGV-Bamani; Mohammad_Mahmudul_static_hgr), super-resolution techniques to enhance image quality (ULTRA-RANGE-UGV-Bamani), attention mechanisms (RAAR3DNet; Benjia_Zhou) to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses (Yunan_Li), and feature selection techniques (SeST) to reduce information redundancy. Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy.
The approaches mentioned previously may be enhanced by using eXplainable Artificial Intelligence (XAI) (Rodis2024Access) techniques, which could be adopted to understand better the behavior of the networks. For instance, Grad-CAM could be further utilized to determine the image region that affects more the predicted gesture class (as has already been done by a few researchers: (CorrNet; SWIN_MSTP_Neurocomputing_2025_CSLR)). Such results may indicate the effectiveness of the embedded on the network attention mechanisms or other related modules to emphasize the most important parts of the input image. t-SNE visualizations could also be leveraged to evaluate the quality of the extracted features, and in particular, how much they are influenced by the presence of redundant information. Thus, guided by the tools provided by the XAI or other similar framework, the interference of gesture-irrelevant factors could be mitigated through the careful design of the future architectures.
@@ 10.2 Computational Cost
Even lightweight methods for SHGR and IDHGR achieve high accuracy. Although, for the more complex task of CHGR, the cost is essential; all RGB-based methods (that are significantly more accurate than pose-based) require more than 366.3 GFLOPs ((DSTEN-CSLR-yu2025dsten)) and reach up to 2950.5 GFLOPs ((TemporalLiftPooling)). This issue prevents not only the deployment of such methods but also the efficient training, which takes longer. Pose-based methods (CoSign) are significantly lighter, but the pose estimation cost is not taken into account. The number of parameters does not change dramatically between IDHGR and CHGR methods, while many SHGR methods require more parameters (e.g. Gesture-CNN (G-CNN-SHARMA2021115657) has 67.25M parameters).
The need of large computational resources does not only prevent the real-time and commercial deployment of VHGR systems, but also put barriers to the evolution of the field, as the conducted experiments take many hours or even days to complete. To mitigate this, authors have focused on reducing the complexity of their models by using non-parametric layers (GestFormer) or designing smaller architectures (LHGR-Net). Thus, reducing the total complexity of the models, could be further examined in the future. Considering also recently proposed approaches for DL models compression (DL_MODEL_COMPRESSION_marino2023deep), such as pruning (i.e., removing less important weights) and quantization (i.e., reducing the arithmetic precision of the parameters), may contribute to this direction. Freezing layers of pretrained models could also be investigated in order to reduce the time required for fine-tuning.
@@ 10.3 Data Scarcity
Many existing datasets either contain a relatively small number of samples or lack sufficient diversity. For example, among static HGR datasets, only the versions of HaGRID (WACV_kapitanov2024hagrid; HAGRIDv2_ARXIV) have been collected from more than one hundred participants, despite some others having a plausible number of samples (LaRED; ULTRA-RANGE-UGV-Bamani). A similar situation is observed in IDHGR datasets, where only the Jester dataset (Jester-Dataset) includes data from over one hundred subjects. Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance (IR_data_imbalance_mobilenet). The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier. This issue is further exacerbated in tasks that involve costly annotation processes and require data-intensive deep learning models with a large number of parameters, such as CSLR.
@@
    Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data (Fangyun_Wei_ICCV).
@@
    Self-Supervised Learning: In the context of the self-supervised learning setting (konstantakos2025self), gloss-to-gloss translation tasks have been used for parameter initialization (CVT-SLR), providing linguistic priors to sign language models. Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure (SignBERT+; SignBERT_ICCV; omar_ikne).
@@
    Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters (ADVERSARIAL_DATA_AUGM), making the models robust to spatial deformations.
@@
    Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed (ZSSLR; FSSLR; ZS-DHGR; ZS-GR; ZERO-SHOT-SLR).
Generative AI, particularly diffusion models (DMs) for image augmentation (alimisis2025advances), could help address data scarcity in HGR. Existing datasets are often limited in size or focused on common tasks, like well-known sign languages. DMs can increase dataset diversity by generating additional samples with varied scenes, subjects, and lighting, or support specialized domains such as rare sign languages and human-robot interaction in firefighting or military scenarios. Thus, leveraging DMs represents a promising direction for future work.
@@ 10.4 Optimization Difficulties
Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR. As shown in CTCA (CTCA), (Fig. 9), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation. Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training (VAC-CSLR; SMKD).
Figure 9: Experimental results in (CTCA) analyze the effect of the temporal aggregation module’s depth—a 1D-TCN with varying layers followed by a two-layer BiLSTM—on training the visual encoder (spatial perception module). In both charts, the x-axis shows the number of 1D-TCN layers. The top chart plots information stored in weights (IIW) (pac-bayes-wang2022pacbayes), indicating generalization ability, while the bottom shows WER. Increasing 1D-TCN layers improves short-term temporal generalization but reduces overall performance due to undertrained visual encoder weights.
@@
    Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision (VAC-CSLR; CSGC_Rao_Sun_Wang_Wang_Zhang_2024; CorrNet; corrnetplus_hu2024corrnetsignlanguagerecognition; TCNet_Lu_Salah_Poppe_2024; C2SLR).
@@
    Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially (CTCA).
@@
    Linguistic Constraints and Gloss Segmentation: Using gloss boundaries (SMKD) or prior knowledge (GPGN) to regularize learning.
The aforementioned approaches —particularly auxiliary supervision (VAC-CSLR)— have served as the foundation for many existing methods. However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques. As a result, optimization remains a significant challenge —one that may be further exacerbated by the continued development of even deeper models.
@@ 10.5 Limitations of Architectures
    Block-fixed models: The existing architectures are based on hand-crafted design, consisting of block of layers with hyperparameters (e.g., number of channels, convolution kernel size) that are defined manually by the researchers. This approach is suboptimal (RAAR3DNet), as several other possible configurations have not been explored.
@@
    Insufficient graph representations of GCNs: The standard GCNs utilized fixed topology graphs, often based on the shape of the hand, to better capture the interactions between the various skeleton joints. However, this configuration has been proved to be suboptimal (TD-GCN), unable to efficiently capture the temporal dynamics.
@@
    Ineffective multimodal fusion: Out of the existing multimodal methods, most of them perform late fusion to incorporate the information of multiple streams (SignBERT_ICCV; SignBERT+; BEST; Weichao-Zhao; MASA; SKIM; SAM_SLR; jiang2021sign-sam-slr-v2). This scheme does not fully leverage the available modalities, as the features extracted by each stream are isolated from the others during the processing on the backbones. Thus, the obtained vector representations do not benefit from the utilization of multiple modalities.
To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures through searching on a large set of configurations (RAAR3DNet), learnable or time-dependent skeleton adjacency matrices that determine implicitly the optimal way to integrate the features of each joint (DSTA_SLR_hu2024dynamic; DSTSA-CGN_cui2025dstsa; ST-PSM_L-PSM; TD-GCN) and multi-level fusion for dual-stream approaches, which enhances the quality of the extracted features by early incorporating information from various sources (CTFB_Hampiholi).
Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap. For instance, to the best of our knowledge, multi-level and mid-level multimodality fusion have been investigated only in the context of dual-stream approaches. Thus, future work could consider the intermediate fusion of three or more streams, e.g. using fusion modules (MMTM; CTFB_Hampiholi) and identifying the optimal way to connect the different streams.

### v3

@@ 8.3 Other
In addition to standard semaphoric gesture datasets, several specialized datasets have been introduced to support research on more specific gesture types. The First-Person Hand Action (FPHA) dataset (55), published in 2018, contains 1,175 video sequences captured from a first-person perspective by six participants. It focuses on isolated manipulative hand gestures across 45 action classes, offering RGB, depth, and skeleton data, and is particularly valuable for studies involving egocentric interaction. Another notable example is the DP Dataset (141), released in 2023, which, to the best of our knowledge, is the only publicly available benchmark specifically designed for pointing gesture recognition. Collected from 33 participants in two indoor environments, the dataset comprises approximately 2.8 million frames and 6,335 annotated gesture instances. It includes detailed annotations such as gesture timings and pointing directions, enabling both binary gesture classification (gesture vs. non-gesture) and directional estimation tasks. Finally, the SocialGesture (28) focuses on the recognition of natural gestures in multi-person environments. It comprises 9,889 in-the-wild videos containing 42,533 instances and provides fine-grained annotations including bounding boxes, relations (for example, a subject is pointing to a target) and Visual Question Answering (VQA) details.
@@ 8.4 Datasets Comparison
After examining the existing datasets for VHGR with significant impact on the current research, in this subsection we aim to determine pros and cons of the most important ones as well as the use cases that they fit. For gesture-based HCI, HaGRID (95) and its versions (145) seem to be a reliable solution, containing an enormous number of samples corresponding to 19 distinct classes (that can be mapped to computer commands). Regarding that these gestures are also static, even simple architectures - like the ones discussed in sections 4 & 7 - can achieve sufficient performance. Furthermore, this dataset could also be used for pretraining, enhancing the backbone’s expression capability. However, its size implies that training takes a long time, which constitutes a deterrent in some cases, compared to other datasets for HCI like OUHANDS (132). For gesture-based HRI, URGR (18) is the only one for UGV guidance and considers various recognition ranges, which may be crucial for real-world scenarios. Even though it contains a large number of samples, the small number of classes may limit the commands given by gestures and, therefore, the functionalities of the entire gesture-based interface. Regarding SLU datasets, for ASL recognition, OpenASL (178) is the largest one, including almost 100K RGB videos (288h in total) collecting from the web. MSASL (94) comprises almost the same number of subjects, but it is much smaller. Furthermore, OpenASL was proposed for CSLR while for MSASL for IDSLR, thus it may be more suitable for real-world use cases, where the temporal boundaries are not known. BSL-1K (11) is also a very large dataset, proposed for British IDSLR and comprising the largest number of samples (280K, 1060h in total). BOBSL (12) was proposed for British CSLR and comprises videos of 1467h in total making it the dataset with the longest duration. Phoenix (49) and Phoenix T (27) remain the most adequate dataset for German sign language, covering a large sign vocabulary. Nevertheless, the small number of samples and the low diversity - due to the limited number of signers and the constrained environment - make essential the collection of a new, larger and more diverse dataset for DGS. For Chinese SLR, NMFs-CSL (80) is significantly smaller than CSLD (202), CSL (87) and NCSL (200) but includes much more signs (1064 instead of less than 500). CSL-Daily (230) remains the only one for Chinese CSLR, covering a large sign vocabulary of 2000 signs in total.
@@ 9.2 Sequence-to-Sequence Metrics
For continuous HGR (CHGR), particularly in Continuous Sign Language Recognition (CSLR), sequence-level metrics are used. Word Error Rate (WER) (136; 2) is the standard metric, measuring the minimum number of substitutions ($S$), deletions ($D$), and insertions ($I$) needed to transform the predicted sequence into the ground truth sequence of length $N$:
@@
Edit Distance Accuracy (50; 204) is a complementary metric representing the ratio of correctly predicted elements in the sequence:
@@ 10.1 Gesture-Irrelevant Factors
The presence of gesture-irrelevant factors (or redundant information) hinders accurate recognition in diverse real-world scenarios. Some of these factors include the following (189; 9; 89; 18; 158; 227):
@@
Aiming to address these challenges, researchers have proposed segmentation models to crop hand regions and remove background interference (229; 36; 171), object detectors for spatial normalization (18; 9), super-resolution techniques to enhance image quality (18), attention mechanisms (227; 228) to enforce the visual encoder to pay attention on the most informative regions, information-theoretic losses (112), and feature selection techniques (191) to reduce information redundancy. Many of these approaches rely on computationally expensive off-the-shelf models, which increase the complexity of the overall HGR methods and set an upper bound to their accuracy.
The approaches mentioned previously may be enhanced by using eXplainable Artificial Intelligence (XAI) (166) techniques, which could be adopted to understand better the behavior of the networks. For instance, Grad-CAM could be further utilized to determine the image region that affects more the predicted gesture class (as has already been done by a few researchers: (84; 15)). Such results may indicate the effectiveness of the embedded on the network attention mechanisms or other related modules to emphasize the most important parts of the input image. t-SNE visualizations could also be leveraged to evaluate the quality of the extracted features, and in particular, how much they are influenced by the presence of redundant information. Thus, guided by the tools provided by the XAI or other similar framework, the interference of gesture-irrelevant factors could be mitigated through the careful design of the future architectures.
@@ 10.2 Computational Cost
Even lightweight methods for SHGR and IDHGR achieve high accuracy. Although, for the more complex task of CHGR, the cost is essential; all RGB-based methods (that are significantly more accurate than pose-based) require more than 366.3 GFLOPs ((214)) and reach up to 2950.5 GFLOPs ((83)). This issue prevents not only the deployment of such methods but also the efficient training, which takes longer. Pose-based methods (92) are significantly lighter, but the pose estimation cost is not taken into account. The number of parameters does not change dramatically between IDHGR and CHGR methods, while many SHGR methods require more parameters (e.g. Gesture-CNN (176) has 67.25M parameters).
The need of large computational resources does not only prevent the real-time and commercial deployment of VHGR systems, but also put barriers to the evolution of the field, as the conducted experiments take many hours or even days to complete. To mitigate this, authors have focused on reducing the complexity of their models by using non-parametric layers (56) or designing smaller architectures (218). Thus, reducing the total complexity of the models, could be further examined in the future. Considering also recently proposed approaches for DL models compression (129), such as pruning (i.e., removing less important weights) and quantization (i.e., reducing the arithmetic precision of the parameters), may contribute to this direction. Freezing layers of pretrained models could also be investigated in order to reduce the time required for fine-tuning.
@@ 10.3 Data Scarcity
Many existing datasets either contain a relatively small number of samples or lack sufficient diversity. For example, among static HGR datasets, only the versions of HaGRID (95; 145) have been collected from more than one hundred participants, despite some others having a plausible number of samples (74; 18). A similar situation is observed in IDHGR datasets, where only the Jester dataset (131) includes data from over one hundred subjects. Furthermore, the amount of samples for specific modalities, such as IR data, is limited compared with the size of RGB-based datasets, an issue known as data imbalance (199). The lack of large-scale, diverse datasets leads to overfitting and limited generalization, which is closely related to the challenges discussed earlier. This issue is further exacerbated in tasks that involve costly annotation processes and require data-intensive deep learning models with a large number of parameters, such as CSLR.
@@
    Multilingual Data Aggregation: Combining datasets (e.g., CSL and DGS) to exploit cross-lingual visual similarities increases the overall amount of training data (207).
@@
    Self-Supervised Learning: In the context of the self-supervised learning setting (102), gloss-to-gloss translation tasks have been used for parameter initialization (226), providing linguistic priors to sign language models. Pretraining with pose reconstruction has also been investigated to incorporate prior knowledge of hand structure (77; 78; 88).
@@
    Advanced Data Augmentation: Adversarial learning has been employed to optimize augmentation hyperparameters (142), making the models robust to spatial deformations.
@@
    Few-/Zero-shot Learning: Approaches using only a small number of samples to correctly recognize gestures—often based on CLIP—have also been proposed (147; 25; 162; 164; 23).
Generative AI, particularly diffusion models (DMs) for image augmentation (13), could help address data scarcity in HGR. Existing datasets are often limited in size or focused on common tasks, like well-known sign languages. DMs can increase dataset diversity by generating additional samples with varied scenes, subjects, and lighting, or support specialized domains such as rare sign languages and human-robot interaction in firefighting or military scenarios. Thus, leveraging DMs represents a promising direction for future work.
@@ 10.4 Optimization Difficulties
Training the visual encoder effectively remains difficult, especially in deep architectures for CSLR. As shown in CTCA (63), (Fig. 9), increasing the depth of the temporal module (e.g., 1D-TCN followed by BiLSTM) improves its generalization but hampers the encoder’s learning due to gradient attenuation in backpropagation. Furthermore, the CTC spike phenomenon—where the model over-focuses on a few keyframes—can cause rapid overfitting and limit encoder training (136; 69).
Figure 9: Experimental results in (63) analyze the effect of the temporal aggregation module’s depth—a 1D-TCN with varying layers followed by a two-layer BiLSTM—on training the visual encoder (spatial perception module). In both charts, the x-axis shows the number of 1D-TCN layers. The top chart plots information stored in weights (IIW) (206), indicating generalization ability, while the bottom shows WER. Increasing 1D-TCN layers improves short-term temporal generalization but reduces overall performance due to undertrained visual encoder weights.
@@
    Auxiliary Supervision: Adding early-stage classifiers to the encoder or short-term module has been proposed as a strategy to guide training by providing additional supervision (136; 161; 84; 82; 124; 234).
@@
    Temporal Module Redesign: Placing the 1D-TCN and BiLSTM in parallel instead of sequentially (63).
@@
    Linguistic Constraints and Gloss Segmentation: Using gloss boundaries (69) or prior knowledge (64) to regularize learning.
The aforementioned approaches —particularly auxiliary supervision (136)— have served as the foundation for many existing methods. However, the increasing complexity of modern deep architectures often exceeds the effectiveness of these techniques. As a result, optimization remains a significant challenge —one that may be further exacerbated by the continued development of even deeper models.
@@ 10.5 Limitations of Architectures
    Block-fixed models: The existing architectures are based on hand-crafted design, consisting of block of layers with hyperparameters (e.g., number of channels, convolution kernel size) that are defined manually by the researchers. This approach is suboptimal (227), as several other possible configurations have not been explored.
@@
    Insufficient graph representations of GCNs: The standard GCNs utilized fixed topology graphs, often based on the shape of the hand, to better capture the interactions between the various skeleton joints. However, this configuration has been proved to be suboptimal (120), unable to efficiently capture the temporal dynamics.
@@
    Ineffective multimodal fusion: Out of the existing multimodal methods, most of them perform late fusion to incorporate the information of multiple streams (78; 77; 224; 225; 223; 115; 91; 90). This scheme does not fully leverage the available modalities, as the features extracted by each stream are isolated from the others during the processing on the backbones. Thus, the obtained vector representations do not benefit from the utilization of multiple modalities.
To address these limitations, researchers have proposed various strategies, including Neural Architecture Search (NAS) to identify optimal model structures through searching on a large set of configurations (227), learnable or time-dependent skeleton adjacency matrices that determine implicitly the optimal way to integrate the features of each joint (86; 32; 30; 120) and multi-level fusion for dual-stream approaches, which enhances the quality of the extracted features by early incorporating information from various sources (66).
Despite these advances, several aspects of deep learning architectures remain underexplored, leaving a substantial research gap. For instance, to the best of our knowledge, multi-level and mid-level multimodality fusion have been investigated only in the context of dual-stream approaches. Thus, future work could consider the intermediate fusion of three or more streams, e.g. using fusion modules (195; 66) and identifying the optimal way to connect the different streams.